OpenStudy (anonymous):

amount $50,000.00; interest rate 17%; comp. annually; due in 12 years. What is the amount must be payed immediately?

OpenStudy (anonymous):

To find the present value of the $50,000, you need to discount it (using the interest rate) back 12 years. The equation to do so will look like this: Present Value * (1+r)^t = Future Value Where: Present Value = What we’re trying to find (I’ll call it X) r = the interest rate (17%) t = time, in years (12 years until maturity) Future Value = The amount paid out in 12 years time ($50,000) Plug these values in and we get: X * (1+0.17)^12 = 50,000 \[X = 50000\div(1.17)^{12}\] X = $7,598.71 (rounded)

OpenStudy (ar43r):

here is another solution. is a bit more complicated and not so accurate, but more flexible. it is based on the formula named "double time" ln(growth factor)*100/time=growth rate ln(50000/x)*100/12=17 ln(50000/x)*100=204 ln(50000)/ln(x)*100=204 (10.8-ln(x))*100=204 1008-100lnx=204 -100ln(x)=-898 lnx=8,98 x=7942 the accuracy of the solution depends strongly on the accuracy of calculation of ln(50000) and e^8,98

OpenStudy (anonymous):

Thanks you guys